OpenStudy (anonymous):

When p equals three, a equals nine, and a varies directly with p, match each value of p with the corresponding value of a. Match Term Definition p equals five A) a equals twelve p equals four B) a equals eighteen p equals two C) a equals six p equals six D) a equals fifteen

OpenStudy (zzr0ck3r):

\(p=a*k\) for some \(k\in \mathbb{Z}\) \(3=9k\) implies \(k=\frac{1}{3}\) so \(p=\frac{a}{3}a\)

OpenStudy (zzr0ck3r):

err that shuold say \(p=\frac{1}{3}a\)

OpenStudy (anonymous):

p equals five D) p equals four A) p equals two C) p equals six B)

OpenStudy (anonymous):

each number is being multiplied by three so that is how you would solve this

OpenStudy (anonymous):

@joquez14 can you help me with one more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thank you Match each set of ordered pairs with their rise over run. Match Term Definition (4, -2), (3, 1) A) -3/7 (4, -2), (-3, 1) B) -1 (-4, -2), (3, 1) C) -3 (-4, 2), (3, 1) D) -1/7

OpenStudy (zzr0ck3r):

and by help you mean just give the answers

OpenStudy (anonymous):

its 3 am im tired! ive been working on this freakin math since 7 am

OpenStudy (zzr0ck3r):

ok

OpenStudy (zzr0ck3r):

the rules dont have a section about giving answers to tired people, maybe I missed it....

OpenStudy (anonymous):

Bye